# Searcher Competition in Block Building

### v1

@@ 3.1 Implementation
In reality, the values that searchers obtain from different blocks is private information to them. However, in a world with passive block producers [^4] We know from [2] that with active block producers implementation of non-trivial solutions is not possible. This in turn resembles previous results from other context where negative results prevail if incomplete information in a two-sided markets is on both sides of the market [12]. In particular, for assignment games [16], which can be re-interpreted as the special case of our model where the validator has additively separable value, there is no mechanism that implements a core-allocation in dominant strategies, for any domain of valuations when there is at least one profile of valuations for which a core allocation that gives positive value to some searcher exists. For a proof of this ”folk theorem” see e.g. [15]., we can implement the extreme point in the core that gives maximal value to searchers in dominant strategies: Consider the case where the block producer is passive i.e. $v_{V}(B)=0$ for each block $B$. Observe that VCG-payments in this problem are defined by
@@
Suppose for reported value functions $(v_{i})_{i\in\mathcal{S}}$, a block $B^{*}$ and payments $(p_{i})_{i\in\mathcal{S}}$ are chosen by the mechanism. By core-stability and Proposition 1, we have $0\leq v_{i}({B}^{*})-p_{i}\leq v(\mathcal{S})-v(\mathcal{S}\setminus\{i\})$ for searcher $i$, or equivalently
@@
which is stricly less. Therefore $i$ gains from misreporting. ∎
@@ 4 Independent Bundles and Competing Searchers
There are examples of core allocations where the validator receives 00.
@@
The allocation in which the validator receives $x_{V}=M$ and searcher $j$ receives $x_{j}=\sum_{i:j\in\text{argmax}_{j\in\mathcal{S}}v_{ij}}(v_{ij}-M_{i})$ is in the core.
@@ 4.1 Stochastic Model
In this section, we analyze the core when searchers find bundles with some probability and success is independent across searchers and opportunities. Let $v_{ij}$ be a binary random variable, which is $1$ with probability $p$ and 00 with probability $1-p$ [^6] Generalizations of the subsequent results to non unit value that can be different for different opportunities are straightforward. The only assumption needed is that searchers conditional on finding the same opportunity generate the same value from it. We could also accommodate heterogeneous value from the same opportunity as long as the noise is sufficiently bounded. Thus, $p$ measures how easy it is to find an arbitrage (bundle).
@@
If $p=\Theta(\frac{1}{n})$ then with positive probability that is constant in $n$ the validator receives 00 in the searcher-optimal core allocation.
@@
The previous results discuss the value distribution for scenarios where the block value is expected to be positive. If probability shrinks faster than $1/n$ as $n$ grows, then with high probability the produced block has value 00. However, conditional on the block having positive value, all value is captured by searchers in the searcher-optimal core allocation.
As remarked in footnote 5, block builders observe around $\sim 1/3$ of submitted bundles clashing. We can use this number to get approximate values of the parameters in our model: we use as $n=125$ which is the number of addresses that have placed at least $2$ bundles in Ethereum blocks within the last $30$ days prior to writing this paper (excluding addresses that have landed only one bundle gives us a crude way to identify addresses that have a high chance of being the main address used by a searcher) according to the website https://libmev.com/. Then
@@
Let $n:=|\mathcal{S}|$. assume the block has capacity to include $(1-\alpha)m$ transactions where $m:=|\mathcal{A}|$ and $11/m<\alpha<$ 1 is a constant. Then, there is a decreasing function $\phi$ such that if $p=\omega(\frac{\phi(\alpha)}{n})$, the validator gets the entire block value with high probability in any core allocation.
@@
$P[Y_{i}<2\text{ for at least }\alpha m\text{ indices}i]$$\leq$ (${m\choose\alpha m}P[Y_{1}<2]^{\alpha m}$
@@
where the last inequality uses the well-known inequality (${a\choose b}\leq\left(\frac{ae}{b}\right)^{b}$ and the previously obtained inequality (5). Choosing $\phi(\alpha)$ to satisfy
@@ 4.2 Empirical validation
We consider bundle data for the time period between September 2023 and February 2024. Bundles generated through MEV-Share make up a significant fraction of the bundles submitted to the Flashbots block builder. We only considered feasible submitted backrun bundles. Figure 3 displays the frequency of different numbers of backrun bundles submitted for the same target transactions. We observe that the distribution is skewed towards 00 in line with our model for choices of the parameters $p$ and $n$ such that the expectation of finding a successful backrun is low, $pn<1$.
In Figure 4, we display the median profit as a function of the number of submitted bundles for the same target transaction. We observe a positive relation between the competition among searchers and the proposer/user profit. This is in line with our theoretical predictions. Zooming in, on the data, we also observe a jump at $1$, see Figure 5.
Running a linear regression, see Table 1, of log median profits on the number of submitted bundles for the same target transaction gives a significant positive relation between profits and the number of submitted bundles.
@@ 5 Conclusion
We have studied MEV extraction in block building as a function of searcher competition and have argued that the core, and in particular the rule that selects the searcher-optimal point within the core are suitable solution concepts that allow us to make robust prediction about value distribution in MEV extraction. We have further identified a dominant-strategy incentive compatible mechanism, giving searchers their marginal value contribution to the realized block, which would be a theoretically appealing payment mechanism for searchers in the case of passive proposers/builders.
The model and solution concept allowed us to to make sense of stylized facts about the Ethereum block building market: validators capture most of the value most of the time, and searcher with unique edge that are less exposed to competition are able to capture significant value. A natural extension of our model for further research would add correlation between different MEV opportunities to our stochastic model. Such an enhanced model would be particularly suitable to study the competition between different block builders and would possibly make theoretical predictions about concentration and value capture in the builder market.

### v2

@@ 3.1 Implementation
In reality, the values that searchers obtain from different blocks is private information to them. However, in a world with passive block producers [^6] We know from [2] that with active block producers implementation of non-trivial solutions is not possible. This in turn resembles previous results from other context where negative results prevail if incomplete information in a two-sided markets is on both sides of the market [12]. In particular, for assignment games [18], which can be reinterpreted as the special case of our model where the validator has additively separable value, there is no mechanism that implements a core-allocation in dominant strategies, for any domain of valuations when there is at least one profile of valuations for which a core allocation that gives positive value to some searcher exists. For a proof of this ”folk theorem” see, e.g. [17]., we can implement the extreme point in the core that gives maximal value to searchers in dominant strategies: Consider the case where the block producer is passive, that is, $v_{V}(B)=0$ for each block $B$. Observe that VCG-payments in this problem are defined by
@@
Suppose that for reported value functions $(v_{i})_{i\in\mathcal{S}}$, a block $B^{*}$ and payments $(p_{i})_{i\in\mathcal{S}}$ are chosen by the mechanism. By core-stability and Proposition 1, we have $0\leq v_{i}({B}^{*})-p_{i}\leq v(\mathcal{S})-v(\mathcal{S}\setminus\{i\})$ for searcher $i$, or equivalently
@@
which is stricly less. Therefore, $i$ gains from misreporting. ∎
@@ 4 Independent Bundles and Competing Searchers
There are examples of core allocations where the validator receives 0.
@@
The allocation in which the validator receives $x_{V}=M$ and searcher $j$ receives
$x_{j}=\sum_{i:j\in\text{argmax}_{j\in\mathcal{S}}v_{ij}}(v_{ij}-M_{i})$
is in the core.
@@ 4.1 Stochastic Model
In this section, we analyze the core when searchers find bundles with some probability and success is independent across searchers and opportunities. The stochastic model assumes binary-valued opportunities. Note the limitation of the setup, as it overlooks the complexity of MEV dynamics, where opportunities are often correlated, and valuations can vary widely. Let $v_{ij}$ be a binary random variable, which is $1$ with probability $p$ and 0 with probability 1p1-p. [^8] Generalizations of the subsequent results to non unit value that can be different for different opportunities are straightforward. The only assumption needed is that searchers conditional on finding the same opportunity generate the same value from it. We could also accommodate heterogeneous value from the same opportunity as long as the noise is sufficiently bounded. Thus, $p$ measures how easy it is to find an arbitrage (bundle).
@@
If $p=\Theta(\frac{1}{n})$ then with positive probability that is constant in $n$ the validator receives 0 in the searcher-optimal core allocation.
@@
The previous results discuss the value distribution for scenarios where the block value is expected to be positive. If probability shrinks faster than $1/n$ as $n$ grows, then with high probability the produced block has value 0. However, conditional on the block having positive value, all value is captured by searchers in the searcher-optimal core allocation.
As remarked in footnote 7, block builders observe around $\sim 1/3$ of submitted bundles clashing. We can use this number to get approximate values of the parameters in our model: we use as $n=125$ which is the number of addresses that have placed at least $2$ bundles in Ethereum blocks within the last $30$ days prior to writing this paper (excluding addresses that have landed only one bundle gives us a crude way to identify addresses that have a high chance of being the main address used by a searcher)  [^9] According to the website https://libmev.com/.. Then
@@
Let $n:=|\mathcal{S}|$. assume the block has capacity to include $(1-\alpha)m$ transactions where $m:=|\mathcal{A}|$ and $1/m<\alpha<1$ is a constant. Then, there is a decreasing function $\phi$ such that if $p=\omega(\frac{\phi(\alpha)}{n})$, the validator gets the entire block value with high probability in any core allocation.
@@
$P[Y_{i}<2\text{ for at least }\alpha m\text{ indices}i]$ $\leq$ ${m\choose\alpha m}P[Y_{1}<2]^{\alpha m}$
@@
where the last inequality uses the well-known inequality ${a\choose b}\leq\left(\frac{ae}{b}\right)^{b}$ and the previously obtained inequality (5). Choosing $\phi(\alpha)$ to satisfy
@@ 5 Empirical validation
We consider bundle data for the time period between September 2023 and February 2024. Bundles generated through MEV-Share make up a significant fraction of the bundles submitted to the Flashbots block builder. We only considered feasible submitted backrun bundles. Figure 3 displays the frequency of different numbers of backrun bundles submitted for the same target transactions. We observe that the distribution is skewed towards 0. This is in line with our model if we choose parameters $p$ and $n$ such that the expectation of finding a successful backrun is low, $pn<1$.
Figure 3: Number of submitted bundles for the same target transaction. Data is grouped into bins of 5, so that the first bar corresponds to the number of transactions receiving 0-4 backruns, the second to the number of transactions receiving 5-9 backruns etc.
In Figure 4, we display the median profit for validators/users as a function of the number of submitted bundles for the same target transaction. We observe a positive relation between the number of searchers targeting the same transaction and the proposer/user profit for that transaction. This is generally in line with our theoretical predictions. We have to make one important caveat though: the value generated by the searcher is not always directly observable in our data set. This, is because the mechanism used by MEV-share is not eliciting the true value of the searchers (in contrast to the VCG mechanisms that we have analyzed above). Instead, the mechanism can be interpreted as a first-price sealed-bid auction for inclusion of the bundle. For back-runs sometimes the value accrued of the searcher can be inferred by his submitted other transactions, but not always and not perfectly. As the auction is first-price, bidding the full true value is not optimal for the searcher, and we are in a situation of partially hidden information. There is also correlation between the value of the opportunity and the number of searchers competing for it, so that higher bids can be partially explained by higher value. Second highest values among submitted bundles is, however, close to the highest value. Thus, the difference in profit going from 1 to 2 bidders can be explained by competition, as most value is extracted by validator/user if there are two or more bundles competing.
Figure 4: Median profit (log-scale) in ETH as a function of number of submitted bundles for the same target transaction.
Our model predicts very sharply that if values can be elicited perfectly, going from one to two competing searchers for the same target transaction, all value from that opportunity would go to the builder/searcher. As there is potentially hidden information, this sharp prediction, does not need to hold in our data however. But we would still expect the profit for the validator/users to to increase sharply going from 1 to 2 bidders, and there to be a diminishing effect of adding additional bidders. Zooming in, on the data, we indeed observe a jump at $1$, see Figure 5. Going from 1 to 2 searchers competing for the same opportunity raises the profit more than any additional searcher, in line with the theory and the previous considerations.
Figure 5: Median profit (log-scale) in ETH as a function of number of submitted bundles for the same target transaction .
Finally, we run a linear regression, see Table 1, of the logarithmic median profits on the number of bundles submitted for the same target transaction. Although the number of observations is not very large, the regression gives a statistically significant positive relation (triple stars mean that the probability hypothesis is wrong is less than 0.001) between profits and the number of submitted bundles. An additional searcher competing for an opportunity raises profits of validator/users by $1.1\%$ on average.
Table 1:
OLS
const -6.087***
received_backrun 0.011***
R-squared 0.340
No. of observations 110
6 Multiple Concurrent Proposers
In this section, we will look at a setting with multiple concurrent proposers in which every next block is built from the input of designated set of proposers. Formally, similar to our approach so far, we can model it with finite set $\mathcal{V}$ of validators, instead of one validator. Let $\mathcal{V}_{-i}$ denote the set of all validators without the validator indexed $i$. We make the mild assumption that the set $\mathcal{V}_{-i}$ can build a block that includes all potential opportunities (bundles), for any $i\in\mathcal{V}$. This assumption can be easily justified if the block is large enough, compared to how much gas opportunities in that block time need, and the total number of validators, $|\mathcal{V}|$, is large enough. Formally:
$v(S\cup\mathcal{V}_{-j})=\max_{B\in\mathcal{B}}(\sum_{i\in S}v_{i}(B)+\sum_{p\in\mathcal{V}_{-j}}v_{p}(B)).$ (6)
Moreover, we assume that proposers themselves do not derive any value from building any block, i.e., $v_{p}(B)=0$ for any $p\in\mathcal{V}$ and $B\in\mathcal{B}$. This is a reasonable assumption when proposers only derive value transferred to them from the searchers.
With only this assumption, we can obtain the following negative result:
Proposition 6.
The core of the game is empty when $|\mathcal{S}|>1$.
Proof.
First, note that no validator can get a positive value in the core allocation. The proof is by contradiction. Suppose that the validator indexed $j$ gets a positive allocation, that is, $x_{j}>0$. Then, there is a deviating subset of all other validators $\mathcal{V}_{-j}$ and all searchers $\mathcal{S}$, in which they get allocated all the value, higher than the original allocation, a contradiction to the core condition that there is no deviation that has a higher total value compared to the total allocation. ∎
However, if there is only one searcher, $|\mathcal{S}|=1$, the core is non-empty.
Proposition 7.
If $|\mathcal{S}|=1$, then the core consists of the allocation in which the unique searcher gets all the value.
Proof.
First, we show that the allocation from the proposition claim is in the core. There is no deviating subset of proposers, since they cannot create any value themselves. Next, note that if any proposer was getting a positive allocation, then the subset consisting of the only searcher and all the other proposers would be deviating and getting more total allocation. ∎
The emptiness of the core for the multiple concurrent builder setting points towards an inherent instability of such a setting. In the presence of multiple proposers we would expect frequent updates and cancellations of searchers’ bids until the end of the block building window, as the market is out-of-equilibrium and re-contracting is always profitable for some players.
@@ 7 Conclusion
We have studied MEV extraction in block building as a function of searcher competition and have argued that the core, and in particular the rule that selects the searcher-optimal point within the core are suitable solution concepts that allow us to make robust prediction about value distribution in MEV extraction. We have further identified a dominant-strategy incentive compatible mechanism, giving searchers their marginal value contribution to the realized block, which would be a theoretically appealing payment mechanism for searchers in the case of passive proposers/builders. The model has been extended to multiple concurrent proposer setting, in which we have obtained that the core is empty.
The model and solution concept allowed us to make sense of stylized facts about the Ethereum block-building market: validators capture most of the value most of the time, and searchers with unique edge that are less exposed to competition are able to capture significant value. A natural extension of our model for further research would add correlation between different MEV opportunities to our stochastic model. Such an enhanced model would be particularly suitable to study the competition between different block builders and would possibly make theoretical predictions about concentration and value capture in the builder market.